Fermi National Laboratory
 

Accelerator Update for October 31, 2007 - November 2, 2007

 

Wednesday October 31

The day shift began with Operations monitoring the Tevatron (TeV) store 5668 (18.83E30) and a Recycler stash (125.9E10), with the Antiproton Source (Pbar) stacking (48.0E10), with the Main Injector (MI), Recycler, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the store, and with beam to the SciBooNE / MiniBooNE experiments.

At 8:38 AM, Operations reported that a Linac Klystron RF station (KRF2) was suffering from coil comparator faults. It reset.

A Pbar expert began an emittance calibration study at 9:33 AM.

A SwitchYard (SY) expert began tuning beam to the Meson M01 absorber at 9:36 AM.

Operations turned off the MI low level RF system at 10:03 AM so an expert could load a new code to drive the RF stations.

TeV experts began preparing at 10:09 AM to adjust the interaction points at CDF and D0.

At 11:01 AM, Operations reported that the MIRF system did not work with the new code. The expert began reverting back to the old code. Using the old code, the system was back up by 11:13 AM, but some problems continued.

Pbar studies ended at 11:17 AM and Pbar resumed stacking.

D0 experts reported at 11:48 AM that their new interaction point was “spot on.” However, CDF reported that its new interaction point was better than its old one, but was still not good.

At noon, Operations reported that the MIRF system still wasn’t working. An expert rebooted the MI dampers (I34 dampers) and within 10 minutes everything was back to normal.

Operations reported at 12:59 PM that they had terminated store 5668.

At 1:47 PM, KRF3 tripped off on a coil comparator fault. It reset.

A SY expert reported at 3:29 PM that he had established beam to MT6 and was done with his tuning and done with the beam.

Operations reported at 4 PM that the Recycler had stashed antiprotons once during the day shift.

Operations put the TeV into shot setup at 4:54 PM.

At 6:25 PM, Operations reported that the P1 transfer line was aborting after every transfer of antiprotons from the Recycler to the TeV.

At 6:55 PM, Operations reported that the TeV suffered a sector B1 quench. The cause was due to orbit changes made at TeV sectors B0 and D0 that were never loaded.

At 8:03 PM, MI suffered the first of two MI power supply loop interlock trips in three minutes. The trips were due to “volts to ground” indications at MI-20 and 30. They reset.

At 9:52 PM, Linac LRF2-5 stations tripped off. They all reset.

Operations put the TeV through a dry squeeze at 10:02 PM.

At 10:15 PM, BRF19 stopped working. Operations bypassed the station and brought BRF7 on its place.

At 10:50 PM, LRF2-5 tripped off again. It reset.

Operations reported at 11:57 PM that the Recycler had stashed once during the evening shift.

Thursday November 1

The midnight shift began with Operations monitoring the stash (192.9E10), with Pbar stacking (23.3E10), with the TeV conducting studies, and with Sci /MiniBooNE taking beam.

At 12:05 AM, KRF1 tripped off due to an LCW flow indication. It reset.

Operations put the TeV into shot setup at 1:13 AM.

At 2:12 AM, Operations reported that KRF2 had tripped off about 12 times since midnight on coil comparator faults.

At 2:14 PM, MIRF9 tripped off. It reset.

At 2:40 AM, the Pbar lithium lens tripped off on an external interlock due to a power supply voltage indication. It was reset.

Operations established store 5674 at 3:49 AM with an initial luminosity of 40.4E30.

At 6:40 AM, Operations reported that KRF2 had tripped off about 12 more times on coil comparator faults.

At 7:56 AM, a vacuum expert reported that the Meson system vacuum in the transfer hall was rapidly degrading. They will need an access to investigate.

Operations reported at 7:57 AM that the Recycler had stashed once during the midnight shift.

At 8:24 AM, Operations reported that the first of two Booster East Anode Power Supply trips in 11 minutes were due to power amplifier LCW trips. An operator reset the supply and contacted an expert.

At 9:30 AM, LRF4 tripped off. It reset.

At 10:07 AM, BRF14 tripped off. It reset. Operations reported that no alarm posted for this event. They contacted an expert.

MI experts began a study at 10:19 AM. Recycler experts monitored its effect on the stash. The study lasted 80 minutes. During the study there was a problem with MIRF10 tripping off.

At 12:10 PM, BRF14 tripped off. It reset.

At 12:38 PM, the BPMs at MI-30 stopped working. A reboot brought them back online.>

Pbar experts began a Pbar Debuncher Stochastic cooling study at 1:41 PM.

At 2:28 PM, LRF3 and 5 tripped off. LRF5 came right back on, but it took about 15 minutes to get LRF3 to come back on.

At 3 PM, the BPMs at MI-30 alarmed on a bad board. An expert investigated and then replaced the board. Later he reported that there was another bad board that needed replacing.

Operations reported at 4:01 PM that the Recycler had stashed once during the day shift.

At 4:25 PM, BRF5 tripped off and would not reset. Operators bypassed the station and brought BRF19 in its place.

A BPM expert reported at 4:36 PM that he had completed his repairs at MI-30.

Operations terminated store 5674 at 4:47 PM.

TeV experts began a proton only study store around 4:52 PM.

At 8:10 PM, the Linac buncher tripped off. It reset.

Operations put the TeV through a wet squeeze at 8:53 PM.

Operations put the TeV into shot setup at 10:06 PM.

At 11:15 PM, the Linac buncher tripped off. It reset.

Operations reported at midnight that the Recycler had stashed once during the evening shift.

Friday November 2

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (102.4E10), with Pbar stacking (17.7E10), and with Sci /MiniBooNE taking beam.

Operations established store 5677 at 12:14 AM with an initial luminosity of 46.26E30.

At 12:28 AM, KRF2 tripped off on a coil comparator fault for the first of six times in 45 minutes. It reset.

At 1:16 AM, Operations reported a tripped yard breaker at the TeV sector F23 service building. The Duty Electrician reset the yard breaker. Operators discovered that the M:V105 power supply had tripped off on a ground fault indication. A power supply expert contacted the operators at F23 and they managed to get the power supply on. Pbar resumed stacking at 2:05 AM.

At 2:19 AM, the Pbar lithium lens tripped off for the first of three times over the next two hours due to an external interlock power supply over voltage indication. Operators reset the lens after each trip.

At 2:45 AM, Operations reported that they had been having more problems with KRF2 and that it had been causing beam position problems in Booster. Operations made some adjustments and talked with an expert who said that the coil power supply would have to be changed out in the morning.

At 3:39 AM, the 750 KeV buncher tripped off. It reset.

Operations reported at 7:55 AM that the Recycler had stashed once during the midnight shift.

The Plans for Friday and the Weekend

The plans for today are to put the TeV into shot setup as soon as the TeV experts complete their studies. The Run Coordinator will make time for the KRF2 coil power supply change and will allow Pbar to conduct a cooling study. Next week there will be a 12 hour or more shutdown to move the low beta magnets at CDF in order to adjust their interaction point and to help the TeV orbit and helix.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/05/2007 -->    email Fermilab